40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 496     Entry time: Sun May 25 19:33:16 2008
Author: rana 
Type: Update 
Category: IOO 
Subject: MC Bad after Re-alignment 
The MC pointing was off and the transmitted power was down after John and Andrey brought it back after the bootfest.

I tried getting it back on Friday but was unsuccessful. Today, after the Phoenix Landing, I got it back to someplace
reasonable, but it still seems to be far off. I will check with Rob before we recenter and of the QPDs.

I had to move all of the MC SUS and also align the beam using the IOO periscope. The attached PDF shows some trends
over the last 80 days. You can see that the drop in MC TRANS is about the same as the drop in PMC TRANS.
Attachment 1: e.pdf  2.917 MB  | Hide | Hide all | Show all
e.pdf
Attachment 2: pmc.png  195 kB  Uploaded Mon May 26 02:33:06 2008  | Show | Hide all | Show all
ELOG V3.1.3-